### v2

# Selection of functional predictors and smooth coefficient estimation for scalar-on-function regression models

Abstract
In the framework of scalar-on-function regression models, in which several functional variables are employed to predict a scalar response, we propose a methodology for selecting relevant functional predictors while simultaneously providing accurate smooth (or more generally regular) estimates of the functional coefficients. We suppose that the functional predictors belong to a real separable Hilbert space, while the functional coefficients belong to a specific subspace of this Hilbert space. Such a subspace can be a Reproducing Kernel Hilbert Space (RKHS) to ensure the desired regularity characteristics, such as smoothness or periodicity, for the coefficient estimates. Our procedure, called SOFIA (Scalar-On-Function Integrated Adaptive Lasso), is based on an adaptive penalized least squares algorithm that leverages functional subgradients to efficiently solve the minimization problem. We demonstrate that the proposed method satisfies an appropriate version of the oracle property adapted to the functional setting, even when the number of predictors exceeds the sample size. SOFIA’s effectiveness in variable selection and coefficient estimation is evaluated through extensive simulation studies and a real-data application to predict GDP growth.
footnotetext: Acknowledgments: We thank Ana Maria Kenney, Michael Morin, Hans-Georg Müller, Matthew Reimherr, Bharath Sriperumbudur and Ruodu Wang, as well as participants at the Second Workshop Mathematics for Artificial Intelligence and Machine Learning (Università Bocconi, 2024), at the 2024 Annual Meeting of the Statistical Society of Canada (Memorial University of Newfoundland), at the Canadian Mathematical Society Summer Meeting 2025 (Québec), at the 2025 International Conference on Statistics and Data Science (Vancouver) and at Université Laval (2025) for useful comments. Hedayat Fathi acknowledges the support of the Faculty of Business Administration, Université Laval, and of the Interuniversity Research Center on Enterprise Network, Logistics and Transportation (CIRRELT). Marzia A. Cremona acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC, grant RGPIN-2020-05657), of the Fonds de recherche du Québec Santé (FRQS, grant 2023-2024-JC-339901), of the Social Sciences and Humanities Research Council (SSHRC, grant 430-2020-00358) and of the Faculty of Business Administration, Université Laval. Federico Severino acknowledges the financial support of the Social Sciences and Humanities Research Council (SSHRC, grant 430-2020-00358). This research was enabled by the computing capability of the Digital Research Alliance of Canada.
Keywords: Variable selection, Scalar-on-function regression, Functional Data Analysis, Reproducing Kernel Hilbert Spaces, Regularization, Oracle property
MSC classification: primary 62R10, 62J05; secondary 46E22.
1 Introduction
High-dimensional regression models are crucial in modern data analysis and have numerous applications in several fields such as genetics, finance, and the social sciences. In the extreme case, some or all of the variables included in the model are functions that vary over a continuum, such as time or space. The use of functional data can result in more accurate data analysis, improved pattern extraction, and enhanced visualization. However, functional data also introduce new challenges in statistical modeling. For example, in functional regression models involving a large number of predictors, selecting the most relevant ones is needed to improve model interpretability and predictive accuracy, as well as to reduce computational complexity (see, e.e,  campos2016; chen2020). Standard references for functional data analysis (FDA) are ramsay2005 and kokoszka2017.
There exist three main types of functional linear regression models: scalar response with functional predictors, functional response with scalar predictors, and functional response with functional predictors. In this study, we focus on the first category, known as scalar-on-function regression, and introduce a new approach to simultaneously perform variable selection and smooth coefficient estimation. In particular, we consider the standardized regression model
$Y_{i}=\sum_{j=1}^{p}\left\langle\beta^{\ast}_{j},X_{ij}\right\rangle_{\mathbb{H}}+\varepsilon_{i},\quad i=1,\cdots,n,$ (1)
where $Y_{i}$ is the scalar response, $X_{ij}$ are the $p$ functional predictors belonging to a real separable Hilbert space $(\mathbb{H},\langle\cdot,\cdot\rangle_{\mathbb{H}})$ (in the simplest case $L^{2}[0,1]$), $\varepsilon_{i}$ are independent scalar random errors with mean 0 and variance $\sigma^{2}$, and $n$ is the sample size. For each $j$, the function $\beta^{\ast}_{j}$ quantifies the effect of the $j^{\text{th}}$ functional predictor on the scalar response. Applications of scalar-on-function regression are widespread, including finance (where curves are time series), medicine, and the natural sciences (ullah2013). Although variable selection methods for classic linear regression models are ubiquitous in the statistical literature (see two recent examples in insolia2022; hanke2024), they have only recently been extended to the functional setting (aneiros2022).
In this paper, we focus on situations where several functional predictors are present but only a small number among them are active (i.e., have $\beta^{\ast}_{j}\neq 0$). Even in classical multiple regression models (where no functional variable is involved), when the number of predictors is large, best subset selection may be computationally infeasible and stepwise selection provides only locally optimal solutions to the variable selection problem. Lasso can be effective in practice (tibshirani1996), but does not satisfy the oracle property (fan2001). The adaptive lasso method (zou2006) adds a data-driven weighting scheme to the penalty term of lasso to improve the accuracy of the estimates and ensure the validity of an oracle property. Under appropriate assumptions, the adaptive lasso is an oracle estimator even when the number of predictors increases with the sample size (huang2008). Following this literature, in this paper, we propose an adaptive lasso method for scalar-on-function regression that simultaneously performs variable selection and coefficient function estimation, ensuring a desired level of regularity for the coefficients. First, we employ a functional version of (non-adaptive) lasso to filter out some variables and obtain initial estimates of the coefficients for the selected predictors. Second, similarly to adaptive lasso, we set weights as reciprocals of the absolute values of the obtained estimates. The primary challenge in the scalar-on-function setting is that the design matrix does not consist of scalar entries but contains elements of a Hilbert space. This difference introduces several computational and theoretical challenges. To avoid redundancy in terminology, we refer to our method as SOFIA-Lasso (Scalar-On-Function Integrated Adaptive Lasso), or simply SOFIA. The term integrated reflects the integration of two norms, enabling the simultaneous selection of variables and regularization of functional coefficients. From an asymptotical point of view, SOFIA enjoys an oracle property even in the case where the number of variables and the number of active predictors increase as the sample size increases. Specifically, we allow the number of predictors to grow faster than the sample size (even at an exponential rate). The challenges of such situations in multiple regression models are illustrated in huang2008.
In our framework, while predictors belong to a general Hilbert space $\mathbb{H}$, we assume that coefficient functions $\beta^{\ast}_{j}$ belong to a specific dense Hilbert subspace $\mathbb{K}$ that incorporates desirable regularity properties such as smoothness or periodicity. This subspace is defined by the eigenfunctions of a positive self-adjoint operator $K$ on $\mathbb{H}$. When a positive-definite kernel induces the operator $K$, the resulting space forms a reproducing kernel Hilbert space (RKHS; see definitions in hsing2015). Hence, we consider a generalization of the RKHS setting which offers greater flexibility. This framework pursues two primary objectives. First, by choosing an appropriate operator $K$, we ensure the desired level of regularity for the coefficient functions $\beta^{\ast}_{j}$. Second, it results in a spectral decomposition that facilitates the approximation of the objective function within finite-dimensional subspaces. The spaces $\mathbb{H}$ and $\mathbb{K}$ share the same orthonormal basis (given by the eigenfunctions of $K$), hence they are defined on the same algebraic structure, but the topology of $\mathbb{K}$ is finer and induces a stronger norm than that of $\mathbb{H}$. This distinction allows us to define the objective function of SOFIA and the following penalized least squares estimator of the coefficients in eq. (1):
$\hat{\bm{\beta}}=\underset{\bm{\beta}\in\mathbb{K}^{p}}{\operatorname{argmin}}~\frac{1}{2n}~\sum_{i=1}^{n}\left(Y_{i}-\sum_{j=1}^{p}\left\langle\beta_{j},X_{ij}\right\rangle_{\mathbb{H}}\right)^{2}+\lambda\sum_{j=1}^{p}\tilde{w}_{j}\left\|\beta_{j}\right\|_{\mathbb{K}},$ (2)
where $\lambda$ is a tuning parameter and $\tilde{w}_{j}$ are adaptive weights designed to penalize smaller coefficient norms more heavily.
A crucial challenge in functional regression is the infinite-dimensional nature of the problem. This issue is particularly pronounced in scalar-on-function regression, where the empirical covariance matrix contains finite-rank operators on $\mathbb{H}$ and cannot be invertible as an operator. To address this issue, a common approach in the literature is the so-called sieve method (wang2016), which consists in projecting the predictors onto a finite-dimensional subspace of $\mathbb{H}$, denoted by $\mathbb{H}^{(m)}$. This subspace is defined by an orthonormal basis, which may be predefined (e.g., splines), data-driven (e.g., functional principal components), or derived from a general basis (roche2023). In the sieve method, the functional entities in the model are approximated over a sequence of nested finite-dimensional subspaces whose union is dense in the underlying Hilbert space. As the sample size increases, the sieve subspace expands, enabling the estimator to approximate the true functional parameters more accurately. This highlights the nonparametric nature of the problem, where consistent estimation is achieved by letting the complexity of the model (the sieve dimension $m$) grow with the sample size rather than fixing a finite number of parameters. Following this method, we derive the required basis from the eigenfunctions of the operator $K$, which allows us to project both functional predictors and coefficient functions onto the subspaces $\mathbb{H}^{(m)}$ and $\mathbb{K}^{(m)}$, respectively, and to obtain the following finite-dimensional approximation of eq. (2) on $\mathbb{K}^{(m)}$:
$\hat{\bm{\beta}}^{(m)}=\underset{\bm{\beta}\in\mathbb{K}^{(m)p}}{\operatorname{argmin}}~~\frac{1}{2n}\sum_{i=1}^{n}~\left(Y_{i}-\sum_{j=1}^{p}\left\langle\beta_{j},X_{ij}^{(m)}\right\rangle_{\mathbb{H}}\right)^{2}+\lambda\sum_{j=1}^{p}\tilde{w}_{j}\left\|\beta_{j}\right\|_{\mathbb{K}}.$ (3)
\@fb@secFB
1.1 Related works
The problem of variable selection for scalar-on-function regression models with multiple functional predictors was first examined by cardot2007ozone, which employed best subset selection in the context of ozone prediction. Although a few methods have been proposed since then, the literature on the subject remains relatively limited. Most of these existing methods are based on a group modeling strategy (aneiros2022) which allows replacing functional variables with a group of scalar variables, and hence reformulating the model as a multiple regression model with grouped predictors. Briefly, each functional variable is expanded using a pre-defined basis system (e.g., B-splines or Fourier basis) and the expansion is then truncated to obtain a finite-dimensional representation of the variable. A grouped variable selection method (see, e.g., yuan2006) is then employed to select groups of predictors, which correspond to functional variables. The first two works proposing this strategy are matsui2011, which use Gaussian basis functions combined with a penalized least squares approach, and gertheiss2013, which employ B-spline basis functions with an adaptive group lasso penalty for variable selection. In the latter, the penalty includes two components (controlled by two different parameters) to regulate the sparsity of the solution and enforce smoothness, respectively. While these methods are interesting from an algorithmic point of view, they do not provide theoretical guarantees on the selection of the true active variables.
Among the methods based on a group modeling strategy that provide theoretical guarantees on variable selection performance, we find fan2015, which use a generic basis and a group lasso penalty, and lian2013 and huang2016 which expand the functional data through functional principal components (FPCA) and use penalized least squares and penalized least absolute deviation, respectively. In particular, fan2015 propose a functional additive regression model that includes linear regression. This study provides a rigorous asymptotic analysis, also when the number of predictors exceeds the sample size. However, the assumption that predictors are known to belong to a Sobolev ellipsoid is quite restrictive. huang2016 propose a similar method employing FPCA to expand functional predictors and a different penalty term. While this approach yields similar analytical results, it is more computationally complex, since it requires estimating a set of basis functions for each functional predictor. Finally, lian2013 provides asymptotic results in the case of fixed number of predictors and of active predictors. More recently, boschi2024 introduced FAStEN, an adaptive elastic net method for variable selection in function-on-function regression, which can be applied also to the scalar-on-function case. This work also expands the functional data using FPCA, and proves an asymptotic oracle property.
All the variable selection methods mentioned above only consider the case of functional data belonging to $L^{2}(T)$, where $T$ is a bounded interval in $\mathbb{R}$. To the best of our knowledge, roche2023 is the only work considering predictors in general Hilbert spaces and proving asymptotic and non-asymptotic properties. In particular, it proposes two adaptations of the group lasso that satisfy sharp oracle inequalities for prediction errors. However, it does not prove results on the selection of the true active variables nor on the coefficient estimations, which we prove for SOFIA.
RKHS, which play an essential role in our study, have received considerable attention over the past two decades in functional linear regression. A seminal contribution is represented by yuan2010, who introduced a smoothness regularization method based on RKHS. Since then, RKHS-based approaches have been employed in various FDA studies for different purposes. For the foundational aspects of RKHS in functional regression models, we refer to kadri2010; recent advancements can be found in wang2022. In the context of RKHS-based linear regression, two approaches can be used for parameter estimation, namely the representer theorem and functional subgradients. The former is the dominant approach in the literature due to its theoretical convenience (see, e.g., yuan2010; shin2016; berrendero2019). However, the methods based on the representer theorem tend to be computationally slow (parodi2018). For this reason, SOFIA is based on functional subgradients (bauschke2011, Chapter 16).
Finally, SOFIA takes inspiration from a series of papers by Reimherr and coauthors on variable selection based on lasso regularization. While these papers explore function-on-scalar regression as opposed to scalar-on-function regression, we borrow from them the definition of functional oracle property and of the Hilbert space $\mathbb{K}$. barber2017 propose FS-Lasso, a method based on the (unweighted) group lasso. Although the oracle property is not satisfied, they give some reasonable bounds for estimation and prediction errors. fan2017 propose an adaptive version of the penalty term in barber2017 and prove the (functional) oracle property of the resulting method. Finally, FLAME (parodi2018) satisfies the oracle property for variable selection, while ensuring desired properties such as smoothness or periodicity for the coefficients by requiring that they belong to an RKHS.
\@fb@secFB
1.2 Our contributions
The general objective of this paper is to introduce a method for variable selection in scalar-on-function regression models that ensures the regularity of the estimated coefficients while satisfying the functional oracle property.
Our first theoretical result concerns the estimator of eq. (3) in the (non-adaptive) case of $\tilde{w}_{j}=1$ and states that, under a suitable restricted eigenvalue (RE) condition, the estimation error admits a sharp bound. Although barber2017 established a similar result, we provide a novel proof for predictors that are finite-dimensional truncations of functional data, i.e. vectors. Our main theorem (i.e., Theorem 2) proves the functional oracle property of SOFIA in the spirit of fan2017. That is, SOFIA correctly identifies the true predictors with probability tending to one as the sample size $n$ increases, under some appropriate assumptions on the finite-dimensional projections. If the tails of the true coefficients after truncation decay rapidly, then the estimation error declines with the nonparametric rate $o_{P}(\tau_{m}/n^{1/2})$, where $\tau_{m}$ is the minimum eigenvalue of the empirical covariance in $\mathbb{K}^{(m)}$. Importantly, this result also holds in the topology of $\mathbb{H}$. This responds to an open question of parodi2018.
Another major contribution is the comparison of the variable selection and estimation performance of multiple recent methods in different simulation scenarios. In particular, we compare SOFIA with the standard group Lasso (fan2015), the semi-adaptive and adaptive group Lasso (gertheiss2013), the method of roche2023, and FAStEN (boschi2024).
Finally, we provide a real-world example for predicting Canadian GDP growth using several economic and financial variables. This application presents several challenges, such as handling unbalanced data and high-frequency variables, which we address by representing the predictors as functional objects.
\@fb@secFB
1.3 Organization of paper
Section 2 introduces the framework and assumptions about the data and coefficients, and provides the mathematical notation and technical conditions needed for the theoretical analysis. Section 3 presents the main theoretical results. Section 4 details the implementation procedure based on functional subgradients. Section 5 presents an extensive simulation study and compares SOFIA’s performance with existing approaches. Section 6 provides a real-data application to the prediction of GDP growth. Finally, Section 7 concludes the work.
2 Mathematical framework
In this section, we describe the general structure of the Hilbert space in which the functional predictors are defined and present the model assumptions, including regularity conditions on the coefficient functions and the structure of the trace-class operator that defines the Hilbert subspace where the coefficients reside. We then provide a structured set of assumptions that ensure the uniform convergence of the estimation procedure.
We assume that the predictors $X_{ij}$ are elements of real separable Hilbert space $(\mathbb{H},\|\cdot\|_{\mathbb{H}})$, where the norm is derived from the inner product $\langle\cdot,\cdot\rangle_{\mathbb{H}}$. Note that $\mathbb{H}$ encompasses a wide range of spaces, including $L^{2}[0,1]$ (used in the simulations of Section 5), $\mathbb{R}^{N}$, and more specialized function spaces commonly used in statistics, such as product spaces, Sobolev spaces, Wiener spaces, and RKHS.
To ensure regularization of the coefficient functions $\beta_{j}^{*}$, we follow the approach of parodi2018. We consider a (compact) trace-class operator, $K:\mathbb{H}\longrightarrow\mathbb{H}$, which is positive definite and self-adjoint. The spectral theorem for compact self-adjoint operators implies that $K$ can be decomposed as $K=\sum_{l=1}^{\infty}\theta_{l}v_{l}\otimes v_{l},$ where $\theta_{1}\geq\theta_{2}\geq\dots>0$ are the ordered eigenvalues and $v_{l}\in\mathbb{H}$ are the corresponding eigenfunctions, which constitute an orthonormal basis for $\mathbb{H}$ (Murphy1990, Theorem 2.4.4). The notation $x\otimes y$ denotes the rank-one operator defined by $(x\otimes y)(h):=\langle y,h\rangle_{\mathbb{H}}x$ for any $x,y,h\in\mathbb{H}$. We define the subspace $\mathbb{K}$ of $\mathbb{H}$ as
$\mathbb{K}:=\left\{h\in\mathbb{H}:~\sum_{l=1}^{+\infty}\frac{\left|\left\langle h,v_{l}\right\rangle_{\mathbb{H}}\right|^{2}}{\theta_{l}}=:\left\|h\right\|_{\mathbb{K}}^{2}<{+\infty}\right\},$ (4)
which is a Hilbert space with norm $\|\cdot\|_{\mathbb{K}}$, $\mathbb{K}$. When $\mathbb{H}=L^{2}[0,1]$, many examples can be found in the literature. In particular, if $K$ is defined as an integral operator associated with a kernel function $k(t,s)$ in a compact domain, then $K$ is a trace-class, positive, and self-adjoint operator. As a result, any RKHS can be viewed as a special case of $\mathbb{K}$. For a detailed discussion on compact and trace-class operators in Hilbert spaces, we refer to Murphy1990, Section 2.4. For an overview of various types of kernel, see berlinet2011, Chapter 7. We assume that the coefficient functions $\beta_{j}^{*}$ belong to $\mathbb{K}$, ensuring the desired level of regularity of the coefficients by choosing an appropriate operator $K$. The objective function of SOFIA, shown in eq. (2), utilizes the $\mathbb{K}$-norm in the penalty term, while the least squares term involves the inner product in $\mathbb{H}$. We study the relationship between $\mathbb{K}$ and $\mathbb{H}$ in Subsection LABEL:app:prop1 of the Supplementary material. Intuitively, we show that $\mathbb{K}$ is the range of the operator $K^{1/2}$, with an appropriate topology, and its norm can be characterized by $\|x\|_{\mathbb{H}}=\|K^{1/2}(x)\|_{\mathbb{K}}$. We also show that $\mathbb{K}$ is dense in $\mathbb{H}$ under the norm of $\mathbb{H}$.
\@fb@secFB
2.1 Model
The following assumption establishes the foundational framework for our functional linear model by specifying the structure of the response, the predictors, and the coefficient functions.
Assumption 1.
Let $Y_{i}$ denote the scalar response for subject $i$, and $X_{i}=(X_{i1},\ldots,X_{ip})\in\mathbb{H}^{p}$ represent the deterministic functional predictors. Without loss of generality, we assume that $Y_{i}$ is centered and the predictors standardized, i.e., $\sum_{i=1}^{n}Y_{i}=0$, $\sum_{i=1}^{n}X_{ij}=0$, and $n^{-1}\sum_{i=1}^{n}\|X_{ij}\|^{2}_{\mathbb{H}}=1$, $j=1,\cdots,p$. We assume that they satisfy the functional linear model of eq. (1) with coefficient functions $\bm{\beta}^{*}=(\beta^{*}_{1},\ldots,\beta^{*}_{p})\in\mathbb{K}^{p}\subset\mathbb{H}^{p}$. The errors $\varepsilon_{i}$ are i.i.d. random variables with mean zero and variance $\sigma^{2}$. In addition, for constants $C_{1},C_{2}>0$, we assume that $P\{|\varepsilon_{i}|>t\}\leq C_{1}\exp(-C_{2}t^{2})$ for all $t\geq 0$ and $i=1,2,\ldots,n$.
Moreover, we assume that only the first $p_{0}<p$ predictors are active, while the remaining $p-p_{0}$ predictors do not contribute to $Y_{i}$. Consequently, their corresponding coefficients are zero.
We note that the sub-Gaussian tail condition on the errors is satisfied, for instance, by Gaussian random variables (boucheron2003, Section 2.3). Importantly, in our model the number of predictors $p$ and the number of active predictors $p_{0}$ are not fixed, but can increase with the sample size $n$. In particular, we require $p_{0}$ to be much smaller than both $n$ and $p$, while we allow $p$ to grow even exponentially with $n$ in our asymptotic results.
\@fb@secFB
2.2 Finite-dimensional approximation
In infinite-dimensional Hilbert spaces, the covariance operator is typically non-invertible due to its finite rank. To address this issue, we restrict the covariance operators to finite-dimensional subspaces of $\mathbb{K}$. In particular, the sieve method (wang2016) enables us to approximate our infinite-dimensional estimator of eq. (2) by solving a sequence of finite-dimensional problems. Following fan2015, we consider a sequence of integers $m=m_{n}=O(n)$, which depends only on the sample size $n$. As $n$ grows to infinity, the dimension $m$ also tends to infinity and the sieve subspaces become dense in $\mathbb{K}$, allowing the estimator to converge to the true coefficient functions. Let $\mathbb{K}^{(m)}$ denote the $m$-dimensional subspace of $\mathbb{K}$ spanned by $\{v_{1},\cdots,v_{m}\}$, and $\mathbb{K}^{(m)\perp}$ its orthogonal complement. The projections of $\beta_{j}\in\mathbb{K}$ on $\mathbb{K}^{(m)}$ and $\mathbb{K}^{(m)\perp}$ are denoted by $\beta_{j}^{(m)}$ and $\beta_{j}^{(m)\perp}$, respectively. We use the analogous notation for the projections of $X_{ij}$ onto $\mathbb{H}^{(m)}$ and $\mathbb{H}^{(m)\perp}$.
In order to approximate the functional coefficients and predictors while controlling the approximation errors, we impose standard assumptions on their bounds.
Assumption 2.
There exist constants $b_{1},b_{2}>0$ and $\gamma>1$ such that $p_{0}=o(m^{(\gamma-1)/2})$, and the active coefficients and predictors satisfy:
$\left\|\beta_{j}^{\ast}\right\|_{\mathbb{K}}\leq b_{1}\quad\text{and}\quad\sup_{1\leq i\leq n}\theta_{l}\left\langle X_{ij},v_{l}\right\rangle_{\mathbb{H}}^{2}\leq b_{2}l^{-\gamma}\quad\text{for all }j=1,\ldots,p_{0}\text{ and }l\geq 1.$
The assumption on the true coefficients implies that such coefficients are uniformly bounded, a standard assumption in infinite-dimensional regression models (fan2015). Since we work with deterministic predictors, the condition on the active predictors is satisfied, for example, when $\|X^{[j]}\|_{\mathbb{H}}$, the column of the design matrix related to the $j^{\text{th}}$ predictor, is uniformly bounded and the operator $K$ exhibits at least a polynomial decay rate ($\theta_{l}\leq Cl^{-\gamma}$ for some constants $C>0$ and $\gamma>1$). The operator $K$ meets this assumption, for example, when it is induced by the Gaussian kernel or the Matérn kernels. Indeed, the Gaussian kernel has exponential decay, while the Matérn kernel has a decay rate of the form $\theta_{l}\leq Cl^{-(1+2\nu)}$ (widom1964), hence the assumption is satisfied for any smoothing parameter $\nu>0$. A condition similar to the bounded design is used in roche2023, where a deterministic design matrix is considered. The condition $p_{0}=o(m^{(\gamma-1)/2})$ reflects the fact that we are working in a sparse regime, where the number of active functional predictors grows more slowly than $m^{(\gamma-1)/2}$, where $m$ is the sieve dimension and $\gamma$ the spectral decay. For kernels with fast spectral decay (e.g., Gaussian kernel or Matérn kernel with $\nu>1$), this requirement is really mild, whereas for kernels with a slower decay rate, such as exponential/Laplacian kernels, it corresponds to the sparsity condition $p_{0}=o(\sqrt{m})$.
The predictors and coefficients can be expanded on the basis $\{v_{l}\}_{l=1}^{+\infty}$ as
$X_{ij}=\sum_{l=1}^{+\infty}\left\langle X_{ij},v_{l}\right\rangle_{\mathbb{H}}v_{l},\qquad\beta_{j}^{\ast}=\sum_{l=1}^{+\infty}\left\langle\beta_{j}^{\ast},v_{l}\right\rangle_{\mathbb{H}}v_{l}.$
If we restrict the predictors and coefficients to the $m$-dimensional subspaces, we obtain the following decomposition for our regression model (1):
$Y_{i}=\sum_{j=1}^{p_{0}}\left\langle\beta_{j}^{*(m)},X_{ij}^{(m)}\right\rangle_{\mathbb{H}}+\sum_{j=1}^{p_{0}}e_{ij}+\varepsilon_{i},$ (5)
where $e_{ij}$ denotes the approximation error, defined as $e_{ij}=\sum_{l=m+1}^{\infty}\langle X_{ij},v_{l}\rangle_{\mathbb{H}}\langle\beta_{j}^{\ast},v_{l}\rangle_{\mathbb{H}}.$ Proposition LABEL:prop:err of Subsection LABEL:app:approx of the Supplementary material shows that, under Assumption 2, these approximation errors converge uniformly to zero as $m\to+\infty$. This permits us to consider the finite-dimensional truncation of model (1):
$Y_{i}=\sum_{j=1}^{p_{0}}\left\langle\beta_{j}^{*(m)},X_{ij}^{(m)}\right\rangle_{\mathbb{H}}+\varepsilon_{i}.$ (6)
To estimate the finite-dimensional coefficients, we consider the objective function of SOFIA:
$L_{\lambda}:\mathbb{K}^{(m)}\longrightarrow\mathbb{R}^{+},\qquad L_{\lambda}(\bm{\beta})=\frac{1}{2n}\sum_{i=1}^{n}\left(Y_{i}-\sum_{j=1}^{p}\left\langle\beta_{j},X_{ij}^{(m)}\right\rangle_{\mathbb{H}}\right)^{2}+\lambda\sum_{j=1}^{p}\tilde{w}_{j}\|\beta_{j}\|_{\mathbb{K}}.$ (7)
The weights $\tilde{w}_{j}$ are chosen adaptively. First, we set all weights to 1 to perform an initial selection and obtain an estimate $\tilde{\beta}_{j}$. Second, we perform an adaptive step with weights $\tilde{w}_{j}=1/\|\tilde{\beta}_{j}\|_{\mathbb{K}}$.
Next, we introduce some notation. We use bold symbols to denote vectors and matrices. Let $\bm{X}=(X_{ij})\in\mathbb{H}^{n\times p}$ be the Hilbert-valued deterministic design matrix, which we partition as $\bm{X}=(\bm{X}_{1},\bm{X}_{2})$, where $\bm{X}_{1}\in\mathbb{H}^{n\times p_{0}}$ and $\bm{X}_{2}\in\mathbb{H}^{n\times(p-p_{0})}$ represent the submatrices of active and null predictors, respectively. The support of $\bm{\beta}^{\ast}$ is defined as $\mathcal{A}=\{1,\ldots,p_{0}\}$ and its complement is $\mathcal{A}^{c}$. When there is no ambiguity, we denote the inner product and norm of both $\mathbb{K}^{p}$ and $\mathbb{K}$ using the same notation.
Since the least squares estimator is based on the inner product of $\mathbb{K}$, we define the covariance operator between two predictors as an operator on $\mathbb{K}$. For $j,k=1,\dots,p_{0}$, the empirical covariance operators are
$\hat{\Gamma}_{jk}:\mathbb{K}\longrightarrow\mathbb{K},\qquad\hat{\Gamma}_{jk}\left(h\right)=\frac{1}{n}\sum_{i=1}^{n}\left\langle h,K\left(X_{ik}\right)\right\rangle_{\mathbb{K}}K\left(X_{ij}\right).$ (8)
These operators incorporate the projection of predictors in the range of $K$. In addition, we define the Hilbert-valued cross-covariance for $j=1,\dots,p_{0}$ as
$\hat{\Gamma}_{YX_{j}}=\frac{1}{n}\sum_{i=1}^{n}Y_{i}K\left(X_{ij}\right)\in\mathbb{K}.$
We denote by $\hat{\Gamma}_{\bm{X_{1}}}$ the $p_{0}\times p_{0}$ matrix of operators $\hat{\Gamma}_{jk}$ and define the vector of cross-covariance operators as $\hat{\Gamma}_{Y\bm{X_{1}}}=(\hat{\Gamma}_{YX_{1}},\dots,\hat{\Gamma}_{YX_{p_{0}}})^{\top}$. The cross-covariance operator-valued matrix between null and active predictors is denoted by $\hat{\Gamma}_{\bm{X_{2}X_{1}}}$ and is defined analogously. For $j,k=1,\cdots,p$, the finite-dimensional covariance operators, indicated by (m), are
$\hat{\Gamma}_{jk}^{(m)}:\mathbb{K}^{(m)}\longrightarrow\mathbb{K}^{(m)},\qquad\hat{\Gamma}_{jk}^{(m)}(h)=\frac{1}{n}\sum_{i=1}^{n}\left\langle h,K\left(X_{ik}^{(m)}\right)\right\rangle_{\mathbb{K}}K\left(X_{ij}^{(m)}\right).$
Finally, we focus on the oracle estimator. This estimator, denoted by $\hat{\bm{\beta}}^{O}=(\hat{\beta}_{1}^{O},\dots,\hat{\beta}_{p}^{O})$, is typically defined as the least squares estimator for active predictors, while it is set to zero for inactive ones. However, in scalar-on-function regression, the empirical covariance matrix $\hat{\Gamma}_{\bm{X_{1}}}$ is always singular, making the least squares estimator inaccessible. In our theory, we work on finite-dimensional spaces $\mathbb{K}^{(m)}$ and define the sieve-oracle estimator as $\bm{\hat{\beta}}_{1}^{O(m)}=(\bm{\hat{\beta}}_{1}^{O(m)},0)$ where $\bm{\hat{\beta}_{1}}^{O(m)}=\hat{\Gamma}_{\bm{X_{1}}}^{(m)-1}\hat{\Gamma}_{Y\bm{X_{1}}}^{(m)}.$ In Section 3, we demonstrate that our estimator satisfies the functional oracle property. Then, we show that under appropriate assumptions, it tends to the true beta at a rate determined by the minimum eigenvalue of the empirical covariance matrix. For more information on the oracle estimator and the functional oracle property, see, for example, fan2017.
\@fb@secFB
2.3 Restricted eigenvalue (RE) condition
In our approach, as a first step, we set all the weights in the objective function (7) to 1. The obtained estimator does not achieve the oracle property, similar to the scalar case (fan2001). Nevertheless, in Theorem 1, we show that under an appropriate version of the RE condition, a tight bound can be found for the distance between the true and the estimated coefficients. We use this bound in the adaptive step to prove the oracle property.
When the predictors are functional, the classic definition of the RE condition is never met: when the dimension of the underlying space is infinity, the lower bound in the RE condition is always zero (roche2023, Lemma 2.1). To overcome this issue, roche2023 proposes to define the RE condition through a sequence ${\alpha}^{(m)}_{n}$ of nonnegative (possibly zero) values. We adopt a similar approach and consider the following condition.
Assumption 3.
(RE condition) For the dimension $m$, there exists a positive number $\alpha_{m}$ such that
$\alpha_{m}:=\min\left\{\frac{\left\|\Gamma_{\bm{XX}}^{(m)1/2}\bm{\delta}\right\|_{\mathbb{K}}}{\left\|\bm{\delta}\right\|_{\mathbb{K}}}:|J|\leq p_{0},~\bm{\delta}=\left(\delta_{1},\dots,\delta_{p}\right)\in\mathbb{K}^{(m)}\setminus\{0\},~\left\|\bm{\delta}_{J^{c}}\right\|_{\ell_{1}/\mathbb{K}}\leq 3\left\|\bm{\delta}_{J}\right\|_{\ell_{1}/\mathbb{K}}\right\},$
where $\|\bm{\delta}\|_{\mathbb{K}}=(\sum_{j=1}^{p}\|\delta_{j}\|_{\mathbb{K}}^{2})^{1/2}$ represents the $\ell_{2}$ norm, $\|\bm{\delta}_{J}\|_{\ell_{1}/\mathbb{K}}=\sum_{j\in J}\|\delta_{j}\|_{\mathbb{K}}$ and $J^{c}$ denotes the complement of the set $J$.
Note that our RE condition differs from that of roche2023 in two key aspects. First, in our definition, we construct the covariance operators using the approximated predictors instead of the predictors. Second, we assume that the dimension $m$ depends on the sample size $n$ and is smaller than it, while roche2023 treats the dimension $m$ and the sample size $n$ as independent, allowing $m$ to grow to infinity while $n$ remains fixed. As a result, in roche2023 for sufficiently large dimensions $\alpha^{(m)}_{n}$ can become zero. In contrast, in our setting $m$ depends on $n$, and $m=O(n)$. In this way, the lower bounds $\alpha^{(m)}$ in Assumption 3 can form a sequence of strictly positive scalars, typically tending to zero. For more details on the RE condition, see bickel2009 for lasso, lounici2011 for group lasso, bellec2017 for a general convex penalty, and barber2017 for function-on-scalar models.
\@fb@secFB
2.4 Asymptotic assumptions
We use the following technical assumptions, which are well-documented in the literature, to assess the asymptotic behavior of our model. If the design matrix is scalar-valued, the upper and lower bounds are typically assumed to be constant (see huang2008; fan2017; parodi2018; mirshani2021). In contrast, following roche2023, we assume that these bounds are sequences that tend to zero as $m$ increases. We study the relation between the RE condition and Assumption 4 in Proposition LABEL:RE in Subsection LABEL:app:RE of the Supplementary material. Specifically, we show that when the covariance between the null predictors tends to zero rapidly, the RE condition is met under Assumption 4.
Assumption 4.
For the dimension $m$, the following conditions hold.
  1. 1.
    Let $\sigma_{\min}$ and $\sigma_{\max}$ be the smallest and largest eigenvalues of $\hat{\Gamma}_{\bm{X_{1}}}^{(m)},$ respectively. There exist a fixed constant $\tau$ and a sequence $\tau_{m}>0$ such that
    $\frac{1}{\tau_{m}}\leq\sigma_{\min}\left(\hat{\Gamma}_{\bm{X_{1}}}^{(m)}\right)\leq\sigma_{\max}\left(\hat{\Gamma}_{\bm{X_{1}}}^{(m)}\right)\leq\tau.$
  2. 2.
    Let $b=\min_{j\in\mathcal{A}}\|\beta_{j}^{\ast}\|_{\mathbb{K}}$$b_{m}=\min_{j\in\mathcal{A}^{(m)}}\|\beta_{j}^{\ast(m)}\|_{\mathbb{K}}$, and $\zeta_{m}=\min\{\tau_{m}^{-1},\alpha_{m}^{2}\}$. Then,
    $\underset{j\in\mathcal{A}}{\max}\|\beta_{j}^{*}-\beta_{j}^{*(m)}\|_{\mathbb{K}}=o(b),\quad\text{and}\quad b_{m}\gg p_{0}\sqrt{\log(p)}/\zeta_{m}\sqrt{n}.$
  3. 3.
    The tuning parameter $\lambda$ satisfies the bounds $p_{0}^{1/2}\log(p)\zeta_{m}^{-1}n^{-1}\ll\lambda\ll b_{m}p_{0}^{-1/2}n^{-1/2}.$
  4. 4.
    Let $\|\cdot\|_{\text{op}}$ denote the operator norm. Then, $\|\hat{\Gamma}_{\bm{X_{2}X_{1}}}^{(m)}\hat{\Gamma}_{\bm{X_{1}}}^{{}^{(m)}-1}\|_{\text{op}}=\varphi_{m}<1.$
These assumptions are counterparts to some standard assumptions in the high-dimensional regression literature. The first one, often called design matrix condition in the scalar design, ensures that the true predictors exhibit regular behavior and are not excessively correlated. This condition is imposed by setting fixed lower and upper bounds for the eigenvalues. In our approach, since $\hat{\Gamma}_{\bm{X_{1}}}$ is a finite-rank operator on an infinite-dimensional Hilbert space $\mathbb{H}^{p_{0}}$, it is never injective, and so 0 is always an eigenvalue. To address this issue, we restrict our analysis to finite-dimensional subspaces of the underlying infinite-dimensional space, making it natural to impose a sequence of bounds that depend on both the dimension and the sample size. Typically, $(\tau_{m})^{-1}$ has zero as an accumulation point, since, as the sample size increases, $\hat{\Gamma}_{\bm{X_{1}}}^{(m)}$ converges to a compact operator (not necessarily of finite rank), and compact operators usually have 0 as a limit point of their eigenvalues (unless they are of finite rank). This implies that the design matrix becomes increasingly ill-conditioned and nearly singular as $n$ grows. This ill-posedness is controlled by the assumptions on the minimum norm of the coefficients and the tuning parameter. In contrast, when the population covariance $\Gamma_{\bm{X_{1}}}$ is bounded, the upper eigenvalue bound can be taken as fixed, similar to scalar cases. The second group of conditions corresponds to the minimum signal condition. It plays two roles in our study. First, it ensures that sieve truncation does not erase the active coefficients. Second, it allows the smallest $\beta_{j}^{*}$ to vary with the sample size, dimension, lower bound of the eigenvalues, and RE bound, while ensuring that it does not become too small. If $(\tau_{m})^{-1}$ or $\alpha_{m}^{2}$ decay too rapidly in relation to sample size, number of predictors, and number of active predictors, then there is no guaranty that the model is able to select predictors with small norms. The third condition prevents the tuning parameter from growing too rapidly or too slowly as $n$ grows. The lower bound prevents overfitting by controlling false positives, while the upper bound ensures that relevant predictors are not excessively penalized, avoiding false negatives. Finally, the fourth condition, known as the irrepresentable condition, ensures that the true and the null predictors do not exhibit a high correlation. In the literature on scalar predictors, the sequence $\varphi_{m}$ is supposed to be a fixed number, usually lower than 1. This requirement has been established as a necessary condition for having the oracle property: see, e.g., (see, e.g., zhao2006, for formal definitions and proofs).
3 Theoretical results
This section presents the theoretical results of the paper. We begin with a concentration inequality that bounds the tail probability of a weighted sum of Hilbert space-valued random variables and serves as a key component in the proofs of Theorems 1 and 2. This result is an extension of Lemma 1 in the supplementary material of huang2008 to the Hilbert space setting. Its proof is given in Subsection LABEL:app:conc of the Supplementary material.
Lemma 1 (Hilbert-valued concentration inequality).
Let $\bm{X}=(X_{ij})\in\mathbb{H}^{n\times p}$ and $\bm{\varepsilon}=(\varepsilon_{1},\cdots,\varepsilon_{n})$ be as in Assumption 1. For each $j=1,\ldots,p$, define $Z_{j}:=n^{-1}\sum_{i=1}^{n}\varepsilon_{i}X_{ij}$. Then, for any $t>0$,
$P\left(\left\|Z_{j}\right\|_{\mathbb{H}}>t\right)\leq\exp\left(-\frac{nt^{2}}{M}\right),$
where $M>0$ is a constant depending only on the constants $C_{1}$ and $C_{2}$ in Assumption 1.
The following theorem provides the bounds for the estimation and prediction errors of the unweighted scalar-on-function lasso. A similar theorem for classic multiple regression can be found in Chapter 6 of buhlmann2011. The function-on-scalar case is proved in barber2017. Our contribution is two-fold. First, we design a concentration inequality for Hilbert-valued design matrices. Then, we develop the estimation and error analysis in a separable Hilbert space. The non-asymptotic and asymptotic convergence rates are nonparametric counterparts of those of the scalar case. This is because our Hilbert-valued concentration inequality mirrors the scalar structure. These rates are crucially dependent on the sequence $\{\alpha_{m}\}_{m\in\mathbb{N}}$ in Assumption 3. This sequence is supposed to be fixed in classic regression models, so it does not have any impact when the sample size increases. However, in the presence of functional predictors, this sequence always tends to zero and its behavior affects the convergence rate of the estimator.
Theorem 1.
Let Assumptions 1-3 hold. Let $\tilde{\bm{\beta}}^{(m)}$ be any minimizer of the lasso function (7) with weights equal to 1. Let $\delta\in(0,1)$ and $\lambda\geq 2\sqrt{n^{-1}\theta_{1}M\log(p/\delta)},$ where $M>0$ is the constant in Lemma 1. Then,
$P\left(\left\|\Gamma_{\bm{XX}}^{(m)1/2}\left(\tilde{\bm{\beta}}^{(m)}-\bm{\beta}^{\ast(m)}\right)\right\|_{\mathbb{K}}\leq\frac{4\lambda\sqrt{p_{0}}}{\alpha_{m}}\right)\geq 1-\delta$
and
$\quad P\left(\left\|\tilde{\bm{\beta}}^{(m)}-\bm{\beta}^{\ast(m)}\right\|_{\ell_{1}/\mathbb{K}}\leq\frac{8\lambda p_{0}}{\alpha_{m}^{2}}\right)\geq 1-\delta.$
The proof of Theorem 1 is provided in Subsection LABEL:app:th_1 of the Supplementary material. Its usefulness is apparent in Corollary 1, which is a direct result of Theorem 1 and makes (unweighted) lasso a very good starting point for the adaptive model. In this corollary, the upper bounds are selected in a way that is consistent with the literature (lounici2011; barber2017; fan2017; parodi2018).
Corollary 1.
Let Assumptions 1-3 hold, and $\lambda$ decays with rate $\lambda\sim\sqrt{\log(p)/p_{0}n}.$ Then, for any minimizer $\tilde{\bm{\beta}}^{(m)}$ of the lasso function (7) with weights equal to 1, for $n\to+\infty$,
$\sup_{j\in\mathcal{A}^{(m)}}\left\|\tilde{\beta}_{j}^{(m)}-\beta_{j}^{\ast(m)}\right\|_{\mathbb{K}}=O_{P}\left(\sqrt{r_{m}}\right)\quad\text{with}\quad r_{m}:=\frac{p_{0}\log p}{\alpha_{m}^{4}n}.$
We are now in a position to state our main result. We show that, when the sample size grows, SOFIA selects the true predictors and converges to the oracle estimator with probability tending to 1. The proof is provided in Subsection LABEL:app:th_2 of the Supplementary material.
Theorem 2 (Functional oracle property).
Let Assumptions 1-4 hold. For $j=1,\cdots,p$ let $\tilde{\beta}_{j}^{(m)}$ be the estimator of Theorem 1 and $\hat{\beta_{j}}^{(m)}$ a minimizer of the function (7) with weights $\tilde{w}_{j}=\|\tilde{\beta_{j}}^{(m)}\|^{-1}_{\mathbb{K}}$. Then, when $n\to+\infty$,
(i) $P\left(\hat{\mathcal{A}}^{(m)}=\mathcal{A}\right)\to 1.$
(ii) $\left\|\hat{\bm{\beta}}^{(m)}-\hat{\bm{\beta}}^{O(m)}\right\|_{\mathbb{K}}=o_{P}(\tau_{m}/\sqrt{n}).$
(iii) If, in addition, $\underset{j\in\mathcal{A}}{\max}\|\beta_{j}^{\ast^{(m)\perp}}\|_{\mathbb{K}}={o(\tau_{m}/\sqrt{p_{0}n})}$, then, as $n\to+\infty$,
$\left\|\hat{\bm{\beta}}^{(m)}-\bm{\beta}^{\ast}\right\|_{\mathbb{K}}=o_{P}(\tau_{m}/\sqrt{n}).$
The first part of the theorem establishes that SOFIA recovers the true support on the $(m)$-dimensional subspace, with probability tending to one. This ensures consistent variable selection in the high-dimensional functional setting. The second part shows that SOFIA is equivalent to the oracle estimator in the finite-dimensional space $\mathbb{K}^{(m)}$. This means that SOFIA is asymptotically equivalent to the oracle estimator on $\mathbb{K}^{(m)}$. The third part claims that, if the tails of the projections of $\beta_{j}^{*}$ in $\mathbb{K}^{(m)}$ decay sufficiently fast, then the finite-dimensional lasso estimator converges to $\bm{\beta}^{\ast}$ in the strong topology of $\mathbb{K}$. As a result, SOFIA achieves full consistency in $\mathbb{K}$, under appropriate conditions, despite operating through finite-dimensional approximations. The proof of Theorem 2, inspired by huang2008 and parodi2018, relies on the functional subgradient of the objective function, together with a concentration inequality for sums of Hilbert-valued random elements. The latter yields high-probability bounds for the stochastic error term. Such bounds are then combined with the adaptive penalization framework to establish both support recovery and estimation consistency.
4 Implementation
To implement SOFIA, we take inspiration from the FLAME method of parodi2018 for the function-on-scalar regression case. To find the minimizers of the objective function (7), we propose a coordinate descent algorithm based on the functional subgradient (bauschke2011, Chapter 16). In Subsection LABEL:app:subg of the Supplementary material we show that, for $j=1,\ldots,p$, the functional subgradient of the function (7) in $\mathbb{K}^{(m)}$ is
$\frac{\partial}{\partial\beta_{j}}L_{\lambda}\left(\bm{\beta}\right)$ $=-\frac{1}{n}\left(\sum_{i=1}^{n}\left(Y_{i}-\sum_{k=1}^{p}\left\langle\beta_{k},K\left(X_{ik}^{(m)}\right)\right\rangle_{\mathbb{K}}\right)K\left(X_{ij}^{(m)}\right)\right)$
[…]
To update the $j^{th}$ coordinate, we set $\check{\beta}_{j}=n^{-1}\sum_{i=1}^{n}(Y_{i}-\sum_{k\neq j}\langle\hat{\beta}_{k,i}^{[t-1]},K(X_{ik}^{(m)})\rangle_{\mathbb{K}})K(X_{ij}^{(m)})$, where $\hat{\beta}_{k,i}^{[t-1]}$ are the estimates obtained in the previous iteration, and fix the other coordinates. Then, we can separate the contribution of the $j^{th}$ predictor and write
$\frac{\partial}{\partial\beta_{j}}L_{\lambda}\left(\bm{\beta}\right)=-\check{\beta}_{j}+\hat{\Gamma}_{jj}^{(m)}\left(\beta_{j}\right)+\lambda\tilde{w}_{j}\begin{cases}\left\|\beta_{j}\right\|_{\mathbb{K}}^{-1}\beta_{j}&\text{if }\beta_{j}\neq 0,\\ \left\{h\in\mathbb{K}^{(m)}:\left\|h\right\|_{\mathbb{K}}\leq 1\right\}&\text{if }\beta_{j}=0.\end{cases}$
Only for implementation, we assume that $\hat{\Gamma}_{jj}^{(m)}$ is equal to the identity matrix. This is equivalent to a well-known group-wise orthonormality condition in the group lasso (roche2023). We then set the last equation to zero to find the minimum of $L_{\lambda}(\bm{\beta})$, obtaining
$\hat{\beta}_{j}=\begin{cases}0&\text{if }\left\|\check{\beta}_{j}\right\|_{\mathbb{K}}\leq\lambda\tilde{w}_{j},\\ \left(1+\left(\lambda\tilde{w}_{j}\right)\left\|\hat{\beta}_{j}\right\|_{\mathbb{K}}^{-1}\right)^{-1}\check{\beta}_{j}&\text{if }\left\|\check{\beta}_{j}\right\|_{\mathbb{K}}>\lambda\tilde{w}_{j}.\end{cases}$ (9)
The first case in eq. (9) provides an upper bound for $\lambda$. If $\lambda\geq(\tilde{w}_{j}n)^{-1}\|\sum_{i=1}^{n}Y_{i}K(X_{ij}^{(m)})\|_{\mathbb{K}}$ for all $j=1,\cdots,p,$ then all predictors are discarded. Therefore, to select the grid of $\lambda$, we start with this value and decrease it in logarithmically spaced intervals at a prespecified rate until we reach the desired number of values for $\lambda$. The second case in eq. (9) enables us to update non-zero coefficients. For $\hat{\beta}_{j}\neq 0$, we have
$\hat{\beta}_{j}=\left(1+\frac{\left(\lambda\tilde{w}_{j}\right)}{\left\|\hat{\beta}_{j}\right\|_{\mathbb{K}}}\right)^{-1}\check{\beta}_{j}.$
This equation does not have a closed form, even for the scalar design. The problem is that the norm of $\hat{\beta}_{j}$ on the right-hand side is unknown. To overcome this issue, we compute the $\mathbb{K}$-norm and apply Parseval’s identity:
$\left\|\hat{\beta}_{j}\right\|_{\mathbb{K}}^{2}=\sum_{l=1}^{m}\left\langle\left(1+\frac{\lambda\tilde{w}_{j}}{\|\hat{\beta}_{j}\|_{\mathbb{K}}}\right)^{-1}\check{\beta}_{j},v_{l}\right\rangle^{2}_{\mathbb{K}}=\left(1+\frac{\lambda\tilde{w}_{j}}{\|\hat{\beta}_{j}\|_{\mathbb{K}}}\right)^{-2}\sum_{l=1}^{m}\left\langle\check{\beta}_{j},v_{l}\right\rangle^{2}_{\mathbb{K}}.$
After some simplifications, we obtain the equation
$1=\frac{\sum_{l=1}^{m}\left\langle\check{\beta}_{j},v_{l}\right\rangle^{2}_{\mathbb{K}}}{\left(\|\hat{\beta}_{j}\|_{\mathbb{K}}+\lambda\tilde{w}_{j}\right)^{2}}.$
We use the NLOPT_LN_COBYLA algorithm from the R package nloptr to solve this nonlinear equation for $\|\hat{\beta}_{j}\|_{\mathbb{K}}$. Then, we use a coordinate descent algorithm to iteratively compute estimates $\hat{\bm{\beta}}^{[t]}$ for $t=1,2,\ldots$ that converge to $\hat{\bm{\beta}}$. The iterative algorithm stops when the improvement in the $\mathbb{K}$-norm of the estimate of $\bm{\beta}^{*}$ reaches a pre-specified threshold. In particular, as stopping criteria, we use a maximum number of iterations of the nonlinear algorithm, a threshold on the improvement of the estimation error, and a kill switch, which means that the process stops when the number of selected variables exceeds a given number. Finally, we employ 5-fold cross-validation to determine the optimal value of $\lambda$.
As explained above, we run the algorithm twice. First, we set all the weights $\tilde{w}_{j}$ to 1 to obtain an initial estimate of the coefficients. Then, for those that have a non-zero norm, we set $\tilde{w}_{j}=1/\|\tilde{\beta}_{j}^{(m)}\|_{\mathbb{K}}$ and run the algorithm a second time to improve the variable selection and refine the estimation.
In the simulation study and real data analysis, we work with the space $\mathbb{H}=L^{2}$ and assume that $K$ is an integral operator associated with the kernel function $k(t,s)$, i.e., $(Kf)(t)=\int k(t,s)f(s)\,ds$. The space $\mathbb{K}$ corresponds to the RKHS induced by the kernel function $k(t,s)$. In the literature, two common examples are the Gaussian kernel and the exponential or Laplacian kernel:
$k_{\infty}\left(t,s\right)=\exp\left(\frac{-\left|t-s\right|^{2}}{\rho}\right),\qquad\qquad k_{1/2}\left(t,s\right)=\exp\left(\frac{-\left|t-s\right|}{\rho}\right),$
where the parameter $\rho$ controls the length scale. A lower value of $\rho$ results in a kernel that decays more rapidly as the distance |ts||t-s| increases, leading to functions that exhibit more rapid variations over short distances. The Gaussian kernel $k_{\infty}\left(t,s\right)$ is infinitely differentiable and is well suited for modeling very smooth data. The exponential or Laplacian kernel $k_{1/2}\left(t,s\right)$ possesses only one continuous derivative, making it more appropriate for modeling rough data. These kernels belong to a broader class known as Matérn kernels, defined by
$k_{\nu}\left(t,s\right)=\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\frac{\sqrt{2\nu}\left|t-s\right|}{\rho}\right)^{\nu}K_{\nu}\left(\frac{\sqrt{2\nu}\left|t-s\right|}{\rho}\right),$
where $\Gamma$ is the Gamma function and $K_{\nu}$ denotes the modified Bessel function of the second kind (genton2001). The predictors are expanded on the eigenfunctions of the operator $K$. To determine the number of basis functions, we use the minimum of $n$ and the smallest number of eigenvalues that satisfy $\sum_{l=1}^{m}\theta_{l}\geq 0.99\sum_{l=1}^{\infty}\theta_{l}$, which mimics the 99% explanation of data variability in FPCA.
The R code implementing SOFIA, as well as an example with simulated data, is provided on the GitHub page https://github.com/HedayatFathi/SOFIA.
5 Simulation study
In this section, we conduct a comprehensive simulation study to evaluate SOFIA’s performance. We consider both the accuracy of variable selection and the predictive performance in several scenarios. We generate datasets under different model complexities and signal-to-noise ratios, and compare SOFIA with existing methods under various specifications.
We take the simulation framework from gertheiss2013 and introduce some modifications in order to obtain scaled design matrices and curve domain $[0,1]$. In particular, we set $\mathbb{H}=L^{2}[0,1]$ and generate the response according to the model
$Y_{i}=\sum_{j=1}^{p_{0}}\int_{0}^{1}\beta_{j}(t)X_{ij}(t)\,dt+\varepsilon_{i},$
where $\varepsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$. The functional predictors are defined as
$X_{ij}\left(t\right)=0.01\left(\sum_{r=1}^{5}\left(a_{ijr}\sin\left(2\pi t(5-a_{ijr})\right)-m_{ijr}\right)+15\right),$
where $i=1,\ldots,n$ denotes the observations, the sample size is fixed to $n=500$, and the interval $[0,1]$ is discretized using a grid of 50 equidistant points. The coefficients $a_{ijr}$ and $m_{ijr}$ are independently sampled from the uniform distributions $U(0,5)$ and $U(0,2\pi)$, respectively.
Figure 1: The coefficient functions $\beta_{j}(t)$ of the active predictors for (a) $p_{0}=5$ and (b) $p_{0}=10$.
The number of predictors $p$ and of active predictors $p_{0}$, indexed by $j$, vary between the different experimental scenarios. In particular, we consider high-dimensional and moderate scales, and very sparse and less sparse regimes for each scale. In the high-dimensional scale, we include $p=700$ functional predictors, which is particularly challenging due to the large number of features relative to the sample size $(n<p)$. Notably, among the competing methods, only FAStEN (boschi2024) can handle this setting. In the high-dimensional scale, the very sparse regime considers five active predictors ($p_{0}=5$), while the less sparse regime includes 10 active predictors ($p_{0}=10$). In the moderate scale, the very sparse regime contains $p=30$ predictors of which $p_{0}=10$ are active, while the less sparse regime has $p=10$ predictors of which $p_{0}=5$ are active. This moderate scale setting enables broader comparisons across different model complexities with five competing methods. The coefficient functions $\beta_{j}$ of the active predictors are simulated based on the Gamma and exponential density functions, as displayed in Figure 1.
For each scenario, we consider five levels of signal-to-noise ratio, namely $\text{SNR}=0.1,0.5,1,10$, and $100$. The variance of the noise term $\varepsilon_{i}$ is set according to the desired SNR level by $\sigma^{2}=var(\bm{Y}^{true})/SNR$. Here, the noise-free scalar response is defined as $Y^{true}_{i}=Y_{i}-\varepsilon_{i}$, and $var(\bm{Y}^{true})$ denotes its variance over $i=1,\ldots,n$.
The performance of the methods is evaluated based on the average number of true positives (TP) and false positives (FP) across 10 replications. In addition, we assess prediction accuracy by computing the root mean squared error (RMSE) on an independently generated test set of size $n_{\text{test}}=100$.
To choose the tuning parameter $\lambda$ in SOFIA, we examine 100 candidate values spaced logarithmically with a ratio of 0.05 (see Section 4), using a 5-fold cross-validation.
\@fb@secFB
5.1 High-dimensional scale
SNR = 1
(a) Very sparse regime ($p_{0}=5$).
(b) Less sparse regime ($p_{0}=10$).
Figure 2: Comparison of the performance of different kernels under varying parameter values for SNR = 1. Subfigures represent the true positives, false positives, and root mean squared error (RMSE) in the very sparse and less sparse regimes, in the high-dimensional scale ($p=700$).
SNR = 10
(a) Very sparse regime ($p_{0}=5$).
(b) Less sparse regime ($p_{0}=10$).
Figure 3: Comparison of the performance of different kernels under varying parameter values for SNR = 10. Subfigures represent the true positives, false positives, and root mean squared error (RMSE) in the very sparse and less sparse regimes, in the high-dimensional scale ($p=700$).
We first compare the performance of SOFIA using different kernel functions (see Section 4) in the high-dimensional scale simulation scenarios, to highlight the impact of the kernel and the choice of the kernel parameter on variable selection and estimation accuracy. In particular, we consider the Gaussian kernel, the exponential kernel, and the Matérn kernels with $\nu=5/2$ and $\nu=3/2$ (see Section 4). For these simulations, we set the kill switch, i.e. the maximum number of selected predictors, to $2p_{0}$.
Figures 2 and 3 display the performance across a range of kernel parameter values for SNR = 1 and 10, respectively, in both regimes. Across all kernels and parameters, SOFIA consistently achieves near-perfect support recovery for both SNR levels. For SNR = 10, all kernels perform similarly well, with perfect support recovery (Figure 3). As expected, in the more noisy case (SNR = 1) sometimes SOFIA misses an active predictor and/or includes a false positive (Figure 2). For both SNR levels, smoother kernels, especially the Gaussian kernel with high kernel parameters, tend to yield slightly lower RMSE. This behavior is expected, since the simulated data involve smooth coefficient functions and predictors. Results for the high-noise cases (SNR = 0.1 and SNR = 0.5) and the near-noise-free case (SNR = 100) are reported in Subsection LABEL:app:extremeSNR of the Supplementary material, and show that all kernels and parameters exhibit comparable performance. Importantly, performance in terms of true and false positive is still pretty good in the case of SNR = 0.5, while it deteriorates markedly for the very noisy case of SNR = 0.1. In Subsection LABEL:app:true_vs_estimate of the Supplementary material, we show the estimated coefficient functions $\hat{\bm{\beta}}$ along with the true $\bm{\beta}^{*}$ in the very sparse regime, for the Gaussian kernel with $\rho=8$ and for the exponential kernel with $\rho=2$. The first kernel serves as a reference and provides smooth estimates of the coefficients, while the second one captures less smooth structures. In both cases, SOFIA successfully recovers the overall shape of the true coefficients. The apparent fluctuation of the estimates is primarily controlled by the kernel parameter $\rho$ rather than by the kernel family. A small $\rho$ leads to more local fluctuations, whereas a larger $\rho$ produces smoother estimates.
Second, we compare the performance of SOFIA (with Gaussian kernel and $\rho=8$) and FAStEN (boschi2024) across different SNR levels and both sparsity levels. For both methods, we use a kill switch of $2p_{0}$ predictors. Note that FAStEN is the only competing method that can handle this high-dimensional scale simulation.
Table 1 shows the results of this simulation. For all SNR levels but SNR = 0.1, both methods identify all or nearly all true active predictors, with very few false positives. Both methods perfectly recover the true support when the noise is low (SNR = 100 and 10). Remarkably, both methods also correctly identify almost all true active predictors, with very few false positives, for SNR = 1 and 0.5. Instead, the case of SNR = 0.1 proves to be extremely challenging for both methodologies. Importantly, SOFIA consistently achieves a lower RMSE than FAStEN across all SNR levels, while maintaining strong support recovery. For SNR = 0.1 and 0.5, SOFIA generally recovers more active predictors, albeit with slightly higher false positive rate. However, this trade-off results in a better predictive accuracy. At higher SNR levels (SNR = 1, 10, and 100), SOFIA continues to produce a lower RMSE. In general, SOFIA demonstrates superior predictive performance while maintaining competitive (often better) variable selection accuracy.
Very sparse ($p_{0}=5$) Less sparse ($p_{0}=10$)
SNR Method TP FP RMSE TP FP RMSE
FAStEN 1.3 0.8 4.42 0.9 0.6 5.72
0.1 SOFIA 2.3 1.4 3.67 2.1 4.8 4.79
FAStEN 4.9 0.9 2.22 7.8 2.4 2.87
0.5 SOFIA 4.7 0 1.72 8.2 1.5 2.32
FAStEN 5 0.1 1.75 9.6 1.8 2.28
1 SOFIA 4.9 0 1.28 9.4 0.3 1.74
FAStEN 5 0 1.21 10 0.1 1.60
10 SOFIA 5 0 0.68 10 0 0.89
FAStEN 5 0 1.14 10 0 1.52
100 SOFIA 5 0 0.60 10 0 0.73
Table 1: Comparison between SOFIA and FAStEN across different SNR levels in the high-dimensional scale ($p=700$). Bold entries highlight the best-performing method for each metric.
\@fb@secFB
5.2 Moderate scale
Since most existing methods are computationally infeasible for a large number of predictors and/or cannot handle cases where the number of predictors exceeds the sample size (such as the high-dimensional case shown above), we consider a moderate scale simulation setting to allow for a broader performance comparison. We compare the performance of SOFIA (with Gaussian kernel and $\rho=8$) with five other methods: the standard non-adaptive group lasso (fan2015), adaptive lasso (Adapt1 and Adapt2; gertheiss2013), the method of roche2023, and FAStEN (boschi2024). To ensure a fair comparison, all methods are run without a kill switch and with their own default configuration parameters.
Very sparse Less sparse
($p=30,~p_{0}=10$) ($p=10,~p_{0}=5$)
SNR Method TP FP RMSE TP FP RMSE
Standard 8.8 10.1 4.96 5 2.7 3.73
Adapt1 9.1 10.4 5.23 5 3.2 3.9
Adapt2 8.8 8.8 5.14 5 2.5 3.77
0.1 Roche 0 0 5.22 0 0 3.86
FAStEN 2.6 0.6 5.77 2.8 0.2 4.50
SOFIA 6.2 2.3 4.73 4.2 0.4 4.01
Standard 10 11.3 2.50 5 2.6 1.72
Adapt1 10 11.5 2.63 5 2.7 1.87
Adapt2 10 8.8 2.57 5 2 1.77
0.5 Roche 0 0 2.72 0 0 2.02
FAStEN 9.6 1.4 2.93 5 0.2 2.3
SOFIA 9.7 0.6 2.28 5 0 2.04
Standard 10 10.8 1.98 5 2.3 1.26
Adapt1 10 12.3 2.10 5 2.5 1.41
Adapt2 10 7.8 2.04 5 1.7 1.31
1 Roche 0 0 2.23 0 0 1.65
FAStEN 10 0.5 2.34 5 0.2 1.85
SOFIA 9.9 0 1.72 5 0 1.57
Standard 10 5.9 1.33 5 0.6 0.59
Adapt1 10 7.4 1.45 5 1.1 0.83
10 Adapt2 10 4.6 1.43 5 0.5 0.66
Roche 9.66 0 1.25 5 0 1.04
FAStEN 10 0.4 1.64 5 0 1.28
SOFIA 10 0 0.90 5 0 0.92
Standard 10 3.4 1.24 5 0.2 0.48
Adapt1 10 5.1 1.38 5 0.5 0.76
100 Adapt2 10 3 1.36 5 0.2 0.54
Roche 10 0 0.57 5 0 1.00
FAStEN 10 0 1.55 5 0 1.20
SOFIA 10 0 0.74 5 0 0.79
Table 2: Comparison of SOFIA with five competitor methods across different SNR levels in the moderate scale. Bold entries highlight the best-performing method for each metric.
Table 2 summarizes the results. In low noise conditions (SNR = 100 and 10), all methods perform generally well. In particular, SOFIA, FAStEN, and Roche consistently achieve high true positive recovery with minimal false positives across both regimes. Although standard and adaptive lasso sometimes achieve better RMSE and always select all true active predictors, their elevated false-positive rates make them less suitable for tasks that require precise variable selection and parsimonious models. In high noise conditions (SNR = 0.5 and 1), SOFIA and FAStEN effectively balance selection and precision, achieving very high true positive rates with relatively low false positives. In contrast, Roche does not select any feature in these settings (hence showing both 0 true and false positive rates). Both standard and adaptive lasso correctly select all true active predictors but also include several false positives, with Adapt2 performing slightly better than Adapt1 and standard lasso in both regimes. Finally, the very noisy case of SNR = 0.1 is challenging for all methods. However, also in this case SOFIA shows good results both in terms of support recovery and of RMSE.
6 Real data analysis
In this section, we apply SOFIA to a real-world economic dataset to identify the most relevant functional predictors to predict future GDP growth in Canada. Our goal is twofold: first, to assess SOFIA’s capacity to select meaningful variables in a high-dimensional real dataset; second, to evaluate the predictive performance of SOFIA compared to classic forecasting methods and classic functional regression. The time-varying nature of economic and financial data makes FDA highly relevant for their analysis. One challenge in these research areas is the high frequency of the data, which can lead to overfitting, noise, and computational complexity. FDA offers robust methods to handle and analyze high-frequency data by smoothing and leveraging their continuous nature (zhang2023). Another challenge is the heterogeneity of the data time frequencies: some variables are reported monthly, while others are reported quarterly or annually (kapetanios2016). In FDA, different measurement frequencies are easily managed because all variables are transformed into curves on a continuous domain and can then be evaluated on the same grid. For recent applications of FSA in economy and finance see, e.g., reimherr2018; severino2022; cremona2023
The prediction of GDP growth is crucial for macroeconomic analysis, as it helps policymakers, financial institutions, and businesses to make decisions. We are interested in cases where multiple macroeconomic predictors are observed and we want to determine the relevant variables to predict GDP growth. We use a dataset from fortin2022, which provides a comprehensive macroeconomic database on Canadian economic indicators. We consider a balanced version of the data that overcomes the issues related to stationarity, missing data, and discontinuities. We use monthly data from January 1981 to December 2019, resulting in 468 monthly observations. This time frame excludes the unpredictable outbreak of the COVID-19 pandemic. The dataset includes many international, national, and provincial economic variables. To ensure meaningful and interpretable results, we restrict our analysis to national-level indicators and select 20 variables that are commonly used in GDP prediction studies. These variables represent various financial sectors and include predictors such as total GDP growth (GDP_new), industrial production GDP growth (IP_new), retail trade GDP (RT_new) growth, public administration GDP growth (PA_new), oil production (OIL_CAN_new), employment rate (EMP_CAN), unemployment rate (UNEMP_CAN), new housing price index (NHOUSE_P_CAN), housing starts units (Hstart_CAN_new), total building permits (build_Total_CAN_new), M3 money supply (M3), household credit (CRED_HOUS), business credit (CRED_BUS), bank rate (BANK_RATE_L), three-month Treasury bill rate (TBILL_3M), total Canada’s official international reserves (RES_TOT), total imports (Imp_BP_new), total exports (Exp_BP_new), exchange rate between the US and the Canadian dollar (USDCAD_new), and the inflation rate measured by the Consumer Price Index (CPI_ALL_CAN). For complete definitions of the variables, see fortin2022.
We construct functional predictors using monthly data by considering 12 consecutive values as a curve. We do not apply additional smoothing, so that the functional representation preserves the original structure of the data. Indeed, since data have been transformed in the original paper to reach stationarity, smoothing would artificially flatten the curves filtering out important information. To generate enough observations, we employ a rolling window method with a one-month shift. Precisely, each functional observation corresponds to a sequence of 12 consecutive months: the first observation consists of a curve obtained by the data from months 1 to 12, the second from months 2 to 13, and so on. Following this procedure, we obtain a dataset that contains 456 observations and 20 predictors. For the scalar response, we consider the GDP growth of the month immediately following the 12-month predictors. GDP growth is measured as the change in GDP from one month to the next, using the difference of logarithmic values. This transformed variable is called GDP_new in the dataset of fortin2022.
Since we want to compare the predictive performance of SOFIA with other methods, we designate the final 36 months of the dataset as the test set for prediction comparison, while the remaining 420 observations are used for fitting the models. We run SOFIA with an exponential kernel and parameter 2, which is well-suited for handling rough data (the functional predictors are not smoothed) and exhibits good robustness to noise. For the selection of $\lambda$, we use a grid of 100 candidate values with a ratio of 0.1, and a rolling-window cross-validation. Specifically, the first 315 observations (75% of the data) are used as the initial training set, while the next 21 observations (5% of the data) are used as the test set. In each subsequent fold, we shift both the training and test sets forward by 21 observations (5% of the data). This procedure results in a 5-fold rolling-window cross-validation framework.
SOFIA selects three functional predictors, namely the total GDP growth (GDP_new), employment rate (EMP_CAN), and unemployment rate (UNEMP_CAN). Figure 4 presents the estimated $\hat{\bm{\beta}}$ coefficients corresponding to the selected predictors. The estimated coefficient functions reveal that the GDP growth and employment rate have a positive relation with the GDP growth of the next month, and their effects grow stronger over time until the 10th month. This suggests that, in the prediction of GDP growth, the GDP growth rate observed three months before has the highest predictive power relative to the other months. In contrast, the unemployment rate shows a negative effect, indicating that higher unemployment is associated with lower GDP growth. Similarly to the other two selected variables, the values of the three months preceding the forecast month are the most influential for this predictor. Interestingly, at the beginning of the period, all three predictors show a pattern opposite to what is observed later. This may suggest a form of mean reversion, where early movements are gradually offset or reversed as the economy adjusts over time.
Figure 4: Estimated coefficients ($\hat{\bm{\beta}}$) for the three selected predictors for GDP growth.
We compare the performance of SOFIA with three classic time series models as benchmarks: the AutoRegressive Integrated Moving Average (ARIMA), the AutoRegressive Integrated Moving Average with Exogenous Regressors (ARIMAX; gu2022), and the naïve model that forecasts the next GDP growth as the last observed value of GDP growth. In addition, we consider a simple scalar-on-function regression model (without penalization) that uses only GDP_new as functional predictor. We estimate ARIMA and ARIMAX using the function auto.arima() from the R package forecast (with default parameters), which automatically selects the optimal parameters based on the corrected Akaike Information Criterion. Then, we give the fitted models as input to the Arima() function to forecast GDP growth for the following month on the test set without reestimating the model. In ARIMAX, we include as exogenous variables all the predictors employed by SOFIA at time t1t-1, ensuring a fair comparison in terms of the variables considered. As for the simple scalar-on-function model, we implement it using a non-penalized version of SOFIA (i.e.,  setting $\lambda=0$). To compare the predictive performance of the methods, we calculate the prediction error as the root mean squared error (RMSE) and the mean absolute value (MAE) over the 36 months in the test set.
Table 3 shows the test RMSE and MAE for the size compared methods. For readability, all RMSE and MAE values in the table are multiplied by 100. The ARIMA model selected by auto.arima() is an ARIMA(4,0,5) with a nonzero mean, while the selected ARIMAX model has ARIMA(3,0,0) errors. We observe that SOFIA achieves the lowest prediction error among the considered methods, slightly outperforming ARIMA in both RMSE and MAE. Although the ARIMAX model shows a marginal improvement in MAE compared to ARIMA, its RMSE is higher, suggesting that the inclusion of all exogenous predictors does not consistently enhance forecast accuracy. The Simple scalar-on-function regression model performs moderately well but is clearly outperformed by SOFIA, as expected. This highlights the benefit of selecting informative predictors through the SOFIA procedure rather than relying on a single functional covariate.
Model ARIMA ARIMAX Naïve Simple SoF SOFIA
RMSE 0.2163 0.2361 0.2625 0.2234 0.2125
MAE 0.1781 0.1742 0.1939 0.1830 0.1651
Table 3: Test RMSE and MAE ($\times$ 100) for the different methods.
7 Conclusions and future work
In this paper, we introduced a novel method, named SOFIA, for variable selection in scalar-on-function regression models, which enables the estimation of functional coefficients with a desired level of smoothness (or, more in general, regularity). To achieve this goal, we require the coefficients to lie in an RKHS-type space that determines the smoothness level, while allowing the predictors to reside in a general Hilbert space. We showed that SOFIA satisfies the functional oracle property, even in scenarios where the number of predictors is much larger than the sample size. In an extensive simulation study, we demonstrated the advantages of our method in variable selection and prediction accuracy with respect to other existing methods.
Our approach adopts a sieve truncation on the eigenbasis of the operator $K$ and solves the penalized least squares with functional subgradients. This strategy yields a sequence of finite-dimensional problems on nested subspaces $\mathbb{K}^{(m)}$ and leads to nonparametric convergence rates that explicitly depend on the sieve level and the complexity of the design. However, selecting the optimal value for the dimension $m$ remains an open problem. Moreover, the representer theorem offers an alternative that enables the study of analytical properties without truncation. This approach is more closely aligned with nonparametric methods and is left for future work.
The literature in the area of functional variable selection is limited, and there is a great space for future research. For instance, we assume that the functional predictors are fully observed and are not contaminated by noise, whereas, in practice, the variables are noisy and observed on a finite grid. Considering the effect of noise and grid size remains an open issue. More broadly, SOFIA is a linear regression method with a lasso penalty. As time-varying data acquisition advances, the need for more complex statistical tools (nonlinear methods or more complex penalties) becomes increasingly more important. This evolution introduces numerous theoretical and computational challenges that push the research agenda in FDA and, in particular, in functional regression models.